### 28 Feb 2026 v1

# arXiv:2603.00704v1[stat.ME] 28 Feb 2026

Abstract.
Two strategies are explored for robustifying classical denoising procedures for the Gaussian sequence model. First, the Hodges and Lehmann (1952) restricted Bayes approach is used to reduce sensitivity to the specification of the initial prior distribution. Second, alternatives to the Gaussian noise assumption are explored. In both cases proposals of Huber (1964) and Mallows (1978) play a crucial role.
Version: . Details on all computations are available on request and will be made available on the web in due course.
1. Introduction
The Gaussian sequence model can be viewed as a compound decision problem with observed $X_{i}\sim\mathcal{N}(\theta_{i},1),\;i=1,\dots,n$. The objective is to estimate the $\theta\in\mathbb{R}^{n}$ subject to quadratic loss. We will denote the standard Gaussian density and cumulative by $\varphi$ and $\Phi$ respectively. Observations are assumed to be exchangeable, so their marginal density is given by,
$f_{G}(x)=\int\varphi(x|\theta)dG(\theta),$
for some mixing distribution $G$. Were $G$ known the optimal (Bayes) decision rule is given by Tweedie’s formula, Efron (2011)
$\hat{\theta}_{i}=\delta^{B}(x_{i})=x_{i}+f_{G}^{\prime}(x_{i})/f_{G}(x_{i}).$
When $G$ is unknown various shrinkage procedures have been proposed, initiated by the fundamental papers of Stein (1956) and Robbins (1956). The extensive literature on Stein shrinkage has offered a rich assortment of practical frequentist and Bayesian procedures for improving upon the naive maximum likelihood estimator, $\delta(x_{i})=x_{i}$ in terms of quadratic loss. Among these procedures more recently the nonparametric maximum likelihood estimator (NPMLE) of Kiefer and Wolfowitz (1956),
$\hat{G}=\mbox{argmax}_{G\in\mathcal{G}}\big\{\sum_{i=1}^{n}\log(f_{G}(x_{i})\big\},$
has been proposed as a plug-in estimator for $G$, Jiang and Zhang (2009), Koenker and Mizera (2014) and Soloff et al. (2024). This $G$-modeling strategy – in the terminology of Efron (2019) – performs well in simulations, e.g. Koenker and Mizera (2014), Gu and Koenker (2016), Koenker and Gu (2026), relative to alternatives that attempt to estimate $f_{G}$ directly or makes a priori assumptions about the form of $G$. However, it is obviously subject to the criticism that the Gaussian assumption on the likelihood is quite strong, and priors are never terribly precise.
In what follows we consider two basic strategies for robustifying empirical Bayes procedures. The first, following a proposal of Hodges and Lehmann (1952), seeks protection from excessive confidence in our initial prior on $G$ by bounding pointwise risk thereby offering a compromise between minimax and Bayes decision rules. The second, acknowledges scepticism about the strictly Gaussian form of $\varphi$, the distribution of the model noise. We find that in accordance with familiar robustness lore that modest modifications of of our initial prior or the Gaussian noise assumption can significantly improve performance of empirical Bayes decision rules while sacrificing only modest performance in the event that the initial prior or the Gaussian noise assumptions are valid.
2. Bayes Risk and Brown’s Identity
Bayes risk in the Gaussian sequence model is,
$r(G,\delta)=\int R(\delta,\theta)dG(\theta),$
with
$R(\delta,\theta)=\mathbb{E}_{\theta}[(\delta(X)-\theta)^{2}]=\int(\delta(x)-\theta)^{2}\varphi(x-\theta)dx.$
Plugging the optimal Bayes rule back into $r(G,\delta)$, we obtain Brown’s identity, Brown (1971):
$r(G,\delta^{B})$ $=\int(x-\theta+\frac{f_{G}^{\prime}(x)}{f_{G}(x)})^{2}\varphi(x-\theta)dxdG(\theta)$
$=1+2\mathbb{E}\Big[\Big(\frac{f_{G}^{\prime}(X)}{f_{G}(X)}\Big)^{\prime}\Big]+\mathbb{E}\Big[\Big(\frac{f_{G}^{\prime}(X)}{f_{G}(X)}\Big)^{2}\Big]$
$=1+2\mathbb{E}\Big[\frac{f_{G}^{{}^{\prime\prime}}(X)}{f_{G}(x)}-\Big(\frac{f_{G}^{\prime}(x)}{f_{G}(x)}\Big)^{2}\Big]+\mathbb{E}\Big[\Big(\frac{f_{G}^{\prime}(X)}{f_{G}(X)}\Big)^{2}\Big]$
$=1-\mathbb{E}\Big[\Big(\frac{f_{G}^{\prime}(X)}{f_{G}(X)}\Big)^{2}\Big]$
$=1-I(\Phi*G).$
The second equality follows from Stein’s lemma, the fourth from the fact that $\int f_{G}^{{}^{\prime\prime}}(x)dx=0$, and the last equality from the definition of Fisher information for distributions with absolutely continuous densities.
It may seem curious that Bayes risk of the optimal empirical Bayes rule for the Gaussian sequence model reduces to Fisher information for a scalar location parameter of the convolution distribution $\Phi*G$. We will exploit the latter connection in the next section to consider least favorable alternative priors for an initial prior in which we lack complete confidence. Such modified priors offer some compromise between strictly Bayesian and minimax procedures. Choosing contamination models by minimizing Fisher information is a classical strategy for choosing alternatives for the univariate Gaussian location and regression problems following Huber (1964). The convolution form of the Fisher information for Bayes risk leads us back to an alternative proposal of Mallows (1978) as well.
A natural objection to many empirical Bayes procedures is that they place unjustified reliance on an initial prior. While such procedures may still perform well with respect to ensemble risk, they may also fail spectacularly for some subpopulations or individuals. This concern underlies the limited translation proposal of Efron and Morris (1971). We will see that bounding minimax risk by modifying an initial prior can often serve to soften the impact of these failings.
3. Restricted Bayes Solutions
In an effort to balance minimax and Bayes solutions, Hodges and Lehmann (1952) proposed solving,
$\min_{\delta}r(G_{0},\delta)\;s.t.\;\max_{\theta}R(\delta,\theta)\leq 1+t,$
for an initial prior $G_{0}$ and some $t>0$. Since $\max_{\theta}R(X,\theta)=1$ corresponding to the worst pointwise risk among all decision rules, achieved by the MLE estimator $\delta(X)=X$, the Hodges and Lehmann modified decision rule is thereby constrained to do uniformly well over the entire parameter space with risk bounded by $1+t$, while minimizing the Bayes risk under prior $G_{0}$, hence called the restricted Bayes rule. They show that this is equivalent to solving,
$\min_{\delta}\;\max_{G\in\mathcal{G}_{\epsilon}(G_{0})}r(G,\delta),$
with $\mathcal{G}_{\epsilon}(G_{0})=\{G=(1-\epsilon)G_{0}+\epsilon H\}$ for some $\epsilon\in(0,1)$ depending upon $t$, where $H$ is an arbitrary distribution for $\theta$. Due to the minimax theorem we can switch the minimization and the maximization, and the resulting optimal rule $\delta^{*}$ is the posterior mean of $\theta$ under the least favorable prior from the class $\mathcal{G}_{\epsilon}(G_{0})$. Berger (1985) comments that “it is very difficult to determine such $\delta^{*}$; furthermore, this ’optimal’ $\delta^{*}$ is usually extremely messy and difficult to work with.” On the contrary, with the aid of modern convex optimization techniques we find them quite tractable and elegant.
Bickel (1983) considers the case with $G_{0}$ having point mass one at zero. Then, by the Brown identity, the Hodges and Lehmann problem is equivalent to solving,
$\max_{G\in\mathcal{G}_{\epsilon}(\delta_{0})}(1-I(\Phi*G))=\min_{G\in\mathcal{G}_{\epsilon}(\delta_{0})}I(\Phi*G)$
This is the problem posed by Mallows (1978) motivated by robustness considerations for time-series problems with additive outliers. Mallows conjectured that the least favorable $G$ would be discrete, supported on the integers with mass declining exponentially. Bickel (1983) reports a modified conjecture of Donoho that relaxes the spacing of the Mallows mass points, but is otherwise similar. Neither conjecture seems to be strictly correct, but numerical computations confirm the nearly exponential decay of the mass. Bickel and Collins (1983) provide a detailed discussion of the discrete nature of the Mallows solutions based on the analyticity of the objective function. Our Appendix A on computation of the Mallows least favorable distribution provides further details. See also Johnstone (1994).
We now consider several special cases of the Hodges and Lehmann restricted Bayes approach. In each case we consider not only the Mallows equivalent form of the Hodges-Lehmann modification, but also a Huber alternative that relaxes the Mallows objective of minimizing the Fisher information over convolutions by minimizing over the entire class of contamination distributions for $X$. Taking the least favorable density and plug into the Tweedie formula gives rise the Huber procedure to estimate $\theta$ for each value of $x$.
The Mallows rule has the obvious advantage that it yields a Bayes decision rule while the corresponding Huber procedure does not. This is particularly evident in the third example of Casella and Strawderman where the unrestricted Bayes rule is minimax, so the Hodges-Lehman’s restriction on point-wise risk is unbinding and the restricted Bayes rule coincides with the unrestricted, but the Huber procedure is inadmissible.
Dirac $G_{0}$:
This is the case considered by Bickel (1983) and anticipated by Mallows (1978). Bickel first simplifies the problem of minimizing $I(\Phi*G)$ by considering the relaxed Huber (1964) problem of minimizing $I(F)$ over $\mathcal{F}=\{F=(1-\epsilon)\Phi+\epsilon W\}$ where $W$ is an arbitrary distribution for $X$. The least favorable distribution $F$ has the well known score function,
$f^{\prime}(x)/f(x)=\begin{cases}x&|x|\leq k\\ k\;\text{sign}(x)&|x|>k.\end{cases}$
Applying Tweedie’s formula gives rise to the hard thresholding Huber rule,
$\delta(x)=x+f^{\prime}(x)/f(x)=\begin{cases}0&|x|\leq k\\ x-k\;\text{sign}(x)&|x|>k\end{cases}$
We contrast this with the numerical solution of the corresponding Mallows problem in Figure 1. The Mallows rule offers a soft thresholding alternative to the piecewise linear Huber rule that oscillates around the Huber rule in the tails.
Figure 1. The figure contrasts the restricted Hodges-Lehmann decision rules based on an initial Dirac prior with mass one at zero: the piecewise linear Huber rule imposes hard thresholding near zero while the Mallows rule allows soft thresholding near zero and oscillates around the Huber rule in the tails. Here $\epsilon=0.4$.
Gaussian $G_{0}$:
This is the case considered by Efron and Morris (1971). The initial prior is Gaussian, $G_{0}\sim\mathcal{N}(0,A)$, so the score function of the least favorable Huber density, which minimizes Fisher information of distributions in the contamination class $\mathcal{F}_{\epsilon}=\{F=(1-\epsilon)N(0,A+1)+\epsilon W\}$, is
$f^{\prime}(x)/f(x)=\begin{cases}\frac{1}{A+1}x&|x|\leq k(A+1)\\ k\;\text{sign}(x)&|x|>k(A+1)\end{cases}$
and Tweedie’s formula yields the piecewise linear “limited translation” rule,
$\delta(x)=x+f^{\prime}(x)/f(x)=\begin{cases}\frac{A}{A+1}x&|x|\leq k(A+1)\\ x-k\;\text{sign}(x)&|x|>k(A+1),\end{cases}$
proposed by Efron and Morris. Figure 2 contrasts the Huber and Mallows forms of the Hodges-Lehmann restricted Bayes rule with the unrestricted Bayes rule. We stick to linear shinkage for values of $x$ in the middle but refrain from shrinking at the two tails. Again, the Mallows rule smooths the Huber rule in the center and oscillates around the Huber rule in the tails. Remarkably, Efron and Morris (1971) in their Appendix Figure B’ already illustrates this behavior of the restricted Mallows rule. Extension of this example to James-Stein forms for $G_{0}$ is straightforward although computation of the associated risk is not as simple as shown by Efron and Morris (1972).
Figure 2. The figure contrasts the Huber and Mallows forms of the restricted Hodges-Lehmann decision rules based on an initial Gaussian prior with variance one: the piecewise linear Huber rule is linear in the center while the Mallows rule is smoother near zero and oscillates around the Huber rule in the tails. Again, $\epsilon=0.4$.
Two Point $G_{0}$ I:
This is the case considered by Casella and Strawderman (1981), the initial prior is $G_{0}=0.5\delta_{-1}+0.5\delta_{1}$ and we restrict the domain of the parameter $\theta$ to the interval [1,1][-1,1]. Casella and Strawderman show that the Bayes rule,
$\delta^{B}(x)=\tanh(x)$
attains maximal pointwise risk of
$R(1,\delta^{B})=\mathbb{E}_{X\sim\mathcal{N}(1,1)}[\tanh(X)-\theta)^{2}]\approx 0.45<1,$
at $\theta=\pm 1$. Consequently, $\delta^{B}$ is minimax and $G_{0}$ is least favorable for all $t(\epsilon)\geq 0$ as asserted in Theorem 3.1 of Casella and Strawderman (1981). In contrast, the Huber modification of the Bayes rule,
$\delta^{H}(x)=\begin{cases}x+k&x<-t\\ \delta^{B}(x)&|x|<t\\ x-k&x>t\end{cases}$
with $k=-(\log f_{0})^{\prime}$ and $t$ solves $\int_{-t}^{t}f_{0}(x)dx+\frac{2f_{0}(t)}{K}=(1-\epsilon)^{-1}$, has strictly greater risk for all $\theta\in\Theta=[-1,1]$.
To see this, let $f_{0}(x)=\frac{1}{2}\varphi(x-1)+\frac{1}{2}\varphi(x+1)$, $g^{H}(x)=\delta^{H}(x)-x$ and $g^{B}(x)=\delta^{B}(x)-x=(\log f_{0}(x))^{\prime}$. By Stein’s lemma,
$R(\theta,\delta^{H})$ $-R(\theta,\delta^{B})$
$=2\mathbb{E}_{\theta}[(g^{H}(X))^{\prime}-(g^{B}(X))^{\prime}]+\mathbb{E}_{\theta}[(g^{H}(X))^{2}-(g^{B}(X))^{2}]$
$=-2\mathbb{E}_{\theta}[1\{|X|>t\}(\log f_{0}(X))^{\prime\prime}]$
$\quad+\mathbb{E}_{\theta}[1\{|X|>t\}(k^{2}-((\log f_{0}(X))^{\prime})^{2}]$
The first term is positive because $f_{0}(x)$ is log-concave. The second term is also positive because $k=-(\log f_{0})^{\prime}(t)$ and $k^{2}>((\log f_{0}(X))^{\prime})^{2}$ for $|x|>t$, hence we conclude $\delta^{H}$ is inadmissible. Figure 3 illustrates the pointwise risk functions of the Bayes, Mallows and Huber decision rules on [1,1][-1,1]. Since the Bayes rule and its Mallows modification are identical the two are indistinguishable in the figure.
Figure 3. Pointwise risk for the Bayes rule and its Huber and Mallows restricted modifications. The Huber risk function assumes $\epsilon=0.4$. The Bayes rule and its Mallows modification are indistinguishable so the Mallows risk has been artificially increased by 0.01 to make them both almost distinguishable.
Two point $G_{0}$ II:
When the two points of support of $G_{0}$ are more widely separated, for example, $G_{0}=\frac{1}{2}\delta_{2}+\frac{1}{2}\delta_{-2}$, and the support of $\theta$ is the whole real line, the marginal $f_{0}(x)=\int\varphi(x-\theta)dG_{0}(\theta)$ is bimodal, the Huber least favaroable density for $X$ is described in the following proposition.
Proposition 3.1.
For $\epsilon$ sufficiently large, [^1] For very small $\epsilon$ no modification in the center of the distribution is required; only the tail behavior is modified. In our example this threshold is about $\epsilon_{0}<0.0001$. the solution to,
$\min_{F\in\mathcal{F}_{\epsilon}}I(F)$
with $\mathcal{F}_{\epsilon}=\{F:F=(1-\epsilon)F_{0}+\epsilon W\}$ where $F_{0}(x)=\frac{1}{2}\Phi(x+2)+\frac{1}{2}\Phi(x-2)$. The least favorable distribution has its density of the form:
$f^{*}(x)=\begin{cases}(1-\epsilon)f_{0}(x)e^{k(x+b)}&x\leq-b\\ (1-\epsilon)f_{0}(x)&|x|\in[c,b]\\ A^{2}\cosh^{2}(kx/2)&x\in[-c,c]\\ (1-\epsilon)f_{0}(x)e^{-k(x-b)}&x\geq b\end{cases}$
where for a given $k$, and $(b,c,A)$,
$k$ $=-(\log f_{0})^{\prime}(b)$
$k\cdot\tanh(kc/2)$ $=(\log f_{0})^{\prime}(c)$
$A^{2}\cosh^{2}(kc/2)$ $=(1-\epsilon)f_{0}(c).$
The constant $k$ is defined implicitly by,
$\int_{-c}^{c}A^{2}\cosh^{2}(kx/2)dx+2\int_{c}^{b}(1-\epsilon)f_{0}(x)dx+\frac{2(1-\epsilon)f_{0}(b)}{k}=1.$
The associated Huber decision rule is then,
$\delta^{*}(x)=\begin{cases}x+k\cdot\tanh(kx/2)&x\in[-c,c]\\ x+f_{0}(x)^{\prime}/f_{0}(x)&|x|\in[c,b]\\ x-k&x\geq b\\ x+k&x\leq-b\end{cases}$
In Figure 4 we compare this shrinkage rule with the initial Bayes rule and the Hodges-Lehmann rule (labeled as Mallows) in the left panel. Again we see that the Mallows’s rule oscillates around the Huber rule in the tails, while being somewhat smoother in the around zero. In the right panel of Figure 4 we depict the pointwise risk of the three procedures for the choice, $\epsilon=0.2$. Worst case risk for the Huber and Mallows rules is about 1.67 while the worst case risk of the Bayes rule is unbounded. The Huber rule is now clearly admissible, but still not as attractive as the Mallows rule, in the sense that its Bayes risk is strictly larger than the Mallows rule.
Figure 4. Decision rules and associated pointwise risk functions for three procedures. The initial prior is: $G_{0}=\frac{1}{2}\delta_{2}+\frac{1}{2}\delta_{-2}$, with contamination level $\epsilon=0.2$. The worst case risk for Mallows and Huber rules is about 1.67.
3.1. Restricted Empirical Bayes rules
Having examined several examples of the Hodges and Lehmann restricted Bayes rules for some simple initial prior distributions, we now consider an empirical Bayes counterpart with $G_{0}$ estimated by maximum likelihood as proposed by Kiefer and Wolfowitz (1956) and anticipated by Robbins (1950). Given a sample from the compound decision problem posed in the introduction, we consider the nonparametric maximum likelihood estimator $\hat{G}$ as the initial $G_{0}$ and then proceed to construct a modified prior according to the principles laid out by Hodges and Lehmann.
Figure 5. The left panel of the figure contrasts the NPMLE prior with the modified Mallows prior. The right panel contrasts the Huber and Mallows forms of the restricted Hodges-Lehmann decision rules based on the initial Kiefer-Wolfowitz NPMLE prior: the Huber rule is almost linear in the center while the Mallows rule is smoother near zero and oscillates around the Huber rule in the tails. In this example we take $\epsilon=0.1$ on the presumption that the initial empirical prior is more reliable than in the prior examples.
We illustrate the consequences of this in Figure 5. Data is generated from the standard Gaussian sequence model with $G_{0}\sim U[0,3]$. Heavy black vertical lines indicate the original NPMLE $\hat{G}$ while the red vertical lines indicate the mass points of the modified prior. While some alteration of the mass in the center of the estimated mixing distribution can be seen, the main change is the new mass points in the tails which decline exponentially in accordance with the Mallows’s conjecture. This feature resembles the Efron and Morris (1971) limited translation estimator that imposed linear shrinkage in the center of the distribution, but eschewed shrinkage in the tails. In baseball terms: a few extremely good hitters deserve their exalted averages. Note that the restricted and unrestricted prior decision rules illustrated in the right panel of the figure agree quite closely on the support of the true $\theta$’s, but diverge sharply beyond this support.
In the following theorem, we establish that the excess Bayes risk, the difference between an oracle Mallows rule $\delta^{M}$ with known $G_{0}$ and an empirical Mallows rule $\hat{\delta}^{M}$ with estimated (NPMLE) $G_{0}$ vanishes asymptotically. The proof makes use of machinery from variational analysis and the important feature that the score function of the oracle and EB Mallows least favorable density is uniformly bounded for the oracle and EB Mallows rule. The proof appears in Appendix C.
Theorem 3.1.
Provided the NPMLE estimator of the marginal density $f_{\hat{G}}$ is Hellinger consistent for the true marginal density $f_{G_{0}}$, then as $n\to\infty$,
$r(G_{0},\hat{\delta}^{M})-r(G_{0},\delta^{M})\to 0$
3.2. Some simulation experience
To evaluate the cost of imposing restrictions on the prior of the Hodges-Lehmann type we consider three examples in this section:
Gaussian $G_{0}$:
$G_{0}\sim\mathcal{N}(0,1)$.
Uniform $G_{0}$:
$G_{0}\sim U[-2,2]$.
Twopoint $G_{0}$:
$G_{0}\sim 0.5(\delta_{-2}+\delta_{2})$.
For each of these settings we compute mean squared error (MSE) for each of the following decision rules:
MLE:
Minimax Rule.
$\delta^{L}$:
Best Linear Rule.
$\delta^{B}$:
Bayes Rule.
$\hat{\delta}_{\hat{G}}^{B}$:
Bayes Rule with NPMLE $\hat{G}$.
$\delta^{H}$:
Huber Modified Bayes Rule.
$\delta^{M}$:
Mallows Modified Bayes Rule.
$\hat{\delta}_{\hat{G}}^{H}$:
Huber Modified Bayes Rule with NPMLE $\hat{G}$.
$\hat{\delta}_{\hat{G}}^{M}$:
Mallows Modified Bayes Rule with NPMLE $\hat{G}$.
The last four rules are evaluated for four distinct values of $\epsilon\in\{0.05,0.1,0.2,0.4\}$. All the simulations are based on 500 replications, for each compound decision problem.
The linear rules work reasonably well for the Gaussian and Uniform settings, however they perform poorly in the two-point setting. The cost of the Hodges-Lehmann restricted priors is modest for small $\epsilon$, but not surprisingly grows substantially when $\epsilon$ is larger. With only $n=100$ observations, the NPMLE rule, $\delta_{\hat{G}}^{B}$, is too variable, but for the larger sample sizes it is nearly competitive with the (oracle) Bayes rules.
MLE $\delta^{L}$ $\delta^{B}$ $\delta_{\hat{G}}^{B}$ $\delta^{H}$ $\delta^{M}$ $\delta_{\hat{G}}^{H}$ $\delta_{\hat{G}}^{M}$
n = 100
$\epsilon=$ 0.05 $0.997$ $0.502$ $0.520$ $0.565$ $0.530$ $0.529$ $0.563$ $0.567$
$\epsilon=$ 0.1 $1.001$ $0.496$ $0.511$ $0.560$ $0.548$ $0.548$ $0.576$ $0.580$
$\epsilon=$ 0.2 $1.006$ $0.504$ $0.520$ $0.566$ $0.604$ $0.600$ $0.622$ $0.622$
$\epsilon=$ 0.4 $0.996$ $0.495$ $0.509$ $0.557$ $0.685$ $0.682$ $0.691$ $0.692$
n = 500
$\epsilon=$ 0.05 $1.001$ $0.501$ $0.503$ $0.518$ $0.529$ $0.529$ $0.533$ $0.534$
$\epsilon=$ 0.1 $1.001$ $0.503$ $0.506$ $0.520$ $0.555$ $0.554$ $0.558$ $0.557$
$\epsilon=$ 0.2 $1.001$ $0.500$ $0.503$ $0.518$ $0.601$ $0.597$ $0.602$ $0.600$
$\epsilon=$ 0.4 $0.999$ $0.501$ $0.504$ $0.519$ $0.689$ $0.686$ $0.688$ $0.686$
n = 1000
$\epsilon=$ 0.05 $0.999$ $0.499$ $0.501$ $0.510$ $0.528$ $0.527$ $0.531$ $0.531$
$\epsilon=$ 0.1 $0.999$ $0.499$ $0.500$ $0.510$ $0.551$ $0.550$ $0.555$ $0.554$
$\epsilon=$ 0.2 $1.006$ $0.502$ $0.504$ $0.513$ $0.603$ $0.599$ $0.603$ $0.601$
$\epsilon=$ 0.4 $0.998$ $0.500$ $0.501$ $0.512$ $0.688$ $0.684$ $0.687$ $0.685$
Table 1. Mean squared error for several compound decision rules. Data generated as $Y=\mu+U$ with $\mu\sim\mathcal{N}(0,1)$ and $U\sim\mathcal{N}(0,1)$, and initial prior is $G_{0}=\mathcal{N}(0,1)$.
MLE $\delta^{L}$ $\delta^{B}$ $\delta_{\hat{G}}^{B}$ $\delta^{H}$ $\delta^{M}$ $\delta_{\hat{G}}^{H}$ $\delta_{\hat{G}}^{M}$
n = 100
$\epsilon=$ 0.05 $1.002$ $0.571$ $0.551$ $0.616$ $0.593$ $0.589$ $0.635$ $0.634$
$\epsilon=$ 0.1 $0.992$ $0.566$ $0.546$ $0.614$ $0.618$ $0.612$ $0.650$ $0.649$
$\epsilon=$ 0.2 $1.006$ $0.573$ $0.553$ $0.620$ $0.678$ $0.671$ $0.700$ $0.699$
$\epsilon=$ 0.4 $1.011$ $0.571$ $0.552$ $0.619$ $0.765$ $0.762$ $0.773$ $0.778$
n = 500
$\epsilon=$ 0.05 $1.004$ $0.570$ $0.550$ $0.568$ $0.592$ $0.588$ $0.598$ $0.595$
$\epsilon=$ 0.1 $1.002$ $0.573$ $0.552$ $0.569$ $0.626$ $0.620$ $0.628$ $0.624$
$\epsilon=$ 0.2 $1.000$ $0.572$ $0.551$ $0.568$ $0.675$ $0.669$ $0.676$ $0.672$
$\epsilon=$ 0.4 $1.004$ $0.573$ $0.554$ $0.572$ $0.762$ $0.760$ $0.761$ $0.760$
n = 1000
$\epsilon=$ 0.05 $0.999$ $0.571$ $0.551$ $0.562$ $0.592$ $0.588$ $0.597$ $0.594$
$\epsilon=$ 0.1 $1.000$ $0.572$ $0.552$ $0.563$ $0.624$ $0.618$ $0.628$ $0.624$
$\epsilon=$ 0.2 $0.999$ $0.570$ $0.550$ $0.561$ $0.674$ $0.668$ $0.676$ $0.671$
$\epsilon=$ 0.4 $1.000$ $0.571$ $0.551$ $0.562$ $0.758$ $0.756$ $0.757$ $0.757$
Table 2. Mean squared error for several compound decision rules. Data generated as $Y=\mu+U$ with $\mu\sim U(-2,2)$ and $U\sim\mathcal{N}(0,1)$, and initial prior is $G_{0}=U(-2,2)$.
MLE $\delta^{L}$ $\delta^{B}$ $\delta_{\hat{G}}^{B}$ $\delta^{H}$ $\delta^{M}$ $\delta_{\hat{G}}^{H}$ $\delta_{\hat{G}}^{M}$
n = 100
$\epsilon=$ 0.05 $1.002$ $0.804$ $0.273$ $0.329$ $0.326$ $0.309$ $0.372$ $0.356$
$\epsilon=$ 0.1 $0.992$ $0.800$ $0.281$ $0.336$ $0.380$ $0.341$ $0.423$ $0.386$
$\epsilon=$ 0.2 $1.006$ $0.809$ $0.285$ $0.347$ $0.490$ $0.416$ $0.531$ $0.466$
$\epsilon=$ 0.4 $1.011$ $0.811$ $0.277$ $0.341$ $0.678$ $0.589$ $0.703$ $0.626$
n = 500
$\epsilon=$ 0.05 $1.004$ $0.803$ $0.280$ $0.294$ $0.329$ $0.316$ $0.345$ $0.330$
$\epsilon=$ 0.1 $1.002$ $0.801$ $0.275$ $0.288$ $0.379$ $0.337$ $0.394$ $0.352$
$\epsilon=$ 0.2 $1.000$ $0.800$ $0.275$ $0.289$ $0.481$ $0.409$ $0.495$ $0.425$
$\epsilon=$ 0.4 $1.004$ $0.805$ $0.279$ $0.292$ $0.672$ $0.583$ $0.682$ $0.599$
n = 1000
$\epsilon=$ 0.05 $0.999$ $0.800$ $0.272$ $0.280$ $0.324$ $0.310$ $0.333$ $0.318$
$\epsilon=$ 0.1 $1.000$ $0.799$ $0.274$ $0.281$ $0.377$ $0.336$ $0.387$ $0.345$
$\epsilon=$ 0.2 $0.999$ $0.799$ $0.273$ $0.281$ $0.480$ $0.408$ $0.490$ $0.418$
$\epsilon=$ 0.4 $1.000$ $0.799$ $0.271$ $0.279$ $0.667$ $0.579$ $0.674$ $0.590$
Table 3. Mean squared error for several compound decision rules. Data generated as $Y=\mu+U$ with $\mu\sim 0.5(\delta_{-2}+\delta_{2})$ and $U\sim\mathcal{N}(0,1)$, and initial prior is $G_{0}=0.5(\delta_{-2}+\delta_{2})$.
4. Robustified Gaussian Likelihoods
Rather than robustifying the prior an alternative strategy is to robustify the likelihood. We will consider two variants of this: the first following Huber (1964) and the second following Mallows (1978).
The classical procedure of Huber for estimating a location parameter is easily adapted to the Gaussian sequence compound decision problem. In place of the Gaussian likelihood in the NPMLE problem we simply insert the Huber log likelihood with density,
$\varphi(u)=\begin{cases}(1-\epsilon)\varphi(k)\exp(-k(u-k))&u>k\\ (1-\epsilon)\varphi(u)&|u|\leq k\\ (1-\epsilon)\varphi(k)\exp(k(u+k))&u<-k\end{cases}$
where $\epsilon$ and $k$ are linked by $2\varphi(k)/k-2\Phi(-k)=\epsilon/(1-\epsilon)$. This density is least favorable, that is has minimal Fisher information for location, in the contamination model,
$\Psi_{\epsilon}=\{\Psi=(1-\epsilon)\Phi+\epsilon H\},$
over all symmetric distributions $H$. When $\epsilon=1/2$ the least favorable Huber distribution is Laplace, or double exponential, and can be viewed as least favorable against asymmetric noise as well as symmetric.
Mallows (1978) proposes to consider minimizing $I(\Phi*G)$ over $\mathcal{G}$, the set of all distributions with mass $1-\epsilon$ at zero, provides an alternative to the Huber contamination model. Rather than assuming iid innovations each arising from the Huber mixture model, Mallows considers an additive outlier model in which with probability $\epsilon$ innovations are standard Gaussian, but occasionally are generated by the convolution $\Phi*H$.
Figure 6. The Huber and Mallows decision rules are contrasted with the Gaussian rule in a setting with $G=U[0,3]$. The heavier tail behavior of the Huber and Mallows base distribution, $\varphi$ results in more aggressive shrinkage with extreme observations discounted as the consequence of noise rather than signal. The Huber and Mallows rules both set $\epsilon=0.1$ for this figure.
In Figure 6 we contrast the Huber and Mallows decision rules with the traditional Gaussian rule. The true mixing distribution, $G$, is chosen to be $U[0,3]$ so the Gaussian rule is itself somewhat curved, not the linear rule we would expect were $G$ itself Gaussian. In contrast the Huber and Mallows rules impose a more aggressive form of shrinkage. With Gaussian $\varphi$ extreme observations can be confidently attributed to signal, while the heavier tailed $\varphi$ of the Huber and Mallows rules tend to attribute such observations to noise. As we have seen previously, the Mallows rule oscillates around the Huber rule in the tails, but otherwise their behavior is quite similar.
When the usual Gaussian $\varphi$ is replaced by either the Huber or Mallows alternative in the nonparametric maximum likelihood estimation of $G$ solutions are accordingly more concentrated with fewer extreme mass points. This effect accentuates the more aggressive shrinkage effect observed in Figure 6.
Both the Mallows and Huber least favorable contamination models offer principled alternatives to the strictly Gaussian noise model. They preserve the convexity of the underlying NPMLE problem and therefore can be easily implemented in software. In the next section we compare performance of several variants of these procedures for a few simulated compound decision settings.
4.1. Some Simulation Experience
We consider the compound decision problem with observations generated from,
$Y_{i}=\theta_{i}+U_{i},\quad i=1,\dots,n,$
with $\theta_{i}$ and $U_{i}$ independent and each generated iidly from $G$ and $\Psi$ respectively. There are two choices of $G$: Either $G\sim U[0,3]$ or $G\sim 0.9\delta_{0}+0.1\delta_{3}$. And three choices of $\Psi$: $\Psi\sim 0.8\Phi+0.2\Phi(\cdot/3)$, $\Psi\sim\text{Laplace}$ and $\Psi=\Phi$, which we label Tukey, Laplace and Gauss respectively.
[…]
Table 4. Mean Squared Error of Several Compound Decision Rules
$G\sim 0.9\delta_{0}+0.1\delta_{3}$ $G\sim U[0,3]$
Tukey Laplace Gauss Tukey Laplace Gauss
Oracle $0.496$ $0.333$ $0.338$ $0.470$ $0.380$ $0.424$
GLmix $0.960$ $0.416$ $0.352$ $1.004$ $0.529$ $0.440$
LLmix $0.688$ $0.398$ $0.457$ $0.631$ $0.476$ $0.482$
HLmix(0.2) $0.696$ $0.584$ $0.658$ $0.659$ $0.708$ $0.672$
HLmix(0.1) $0.693$ $0.400$ $0.429$ $0.689$ $0.561$ $0.521$
HLmix(0.05) $0.779$ $0.364$ $0.371$ $0.793$ $0.495$ $0.463$
HLmix(0.025) $0.857$ $0.370$ $0.354$ $0.885$ $0.487$ $0.443$
MLmix(0.2) $0.674$ $0.567$ $0.636$ $0.612$ $0.676$ $0.613$
MLmix(0.1) $0.664$ $0.390$ $0.421$ $0.654$ $0.545$ $0.502$
MLmix(0.05) $0.751$ $0.357$ $0.367$ $0.759$ $0.485$ $0.458$
MLmix(0.025) $0.837$ $0.363$ $0.352$ $0.861$ $0.477$ $0.441$
It is evident from the table that the Gaussian NPMLE, GLmix, performs best when the noise distribution is actually Gaussian; however, when $\Psi\neq\Phi$ it pays to consider one of the alternatives. The Mallows NPMLE procedures seem to perform slightly better than the corresponding Huber methods, while the Laplace NPMLE, LLmix, which can be regarded as a “median-type” estimator performs surprisingly well over all the experimental settings.
5. conclusion
We have considered two distinct strategies for robustifying empirical Bayes decision rules for the Gaussian sequence model. In the first motivated by the seminal paper of Hodges and Lehmann we would like protection against deviations from an initial Bayes prior. In the second we seek protection against non-Gaussian behavior in the noise distribution. Both strategies rely on the classical robustness proposals of Huber (1964) and Mallows (1978). Some combination of the two strategies is obviously possible, but choice of tuning parameters remains a delicate issue.
Appendix A Computation of the Mallows least favorable distribution
As noted by Donoho and Reeves (2013) the Mallows (1978) problem is convex and therefore is easily solved numerically with modern convex optimization software. Our objective is to solve
$\underset{f\in\mathcal{K}}{\min}\int\frac{f^{\prime}(x)^{2}}{f(x)}dx$
with
$\mathcal{K}_{\epsilon}=\Big\{f=(1-\epsilon)\int\varphi(x-\theta)d\delta_{0}(\theta)+\epsilon\int\varphi(x-\theta)dH(\theta)\Big\}$
This can be solved, on a grid of $x$, $\{x_{1}<x_{2}<\dots<x_{M}\}$ and a grid of $\theta$ as $\{\theta_{1}<\theta_{2}<\dots<\theta_{L}\}$ as the rotated quadratic cone convex optimization problem:
$\min\sum_{i=1}^{M}w_{i}$
subject to
$u_{i}$ $=f_{i+1}-f_{i},$
$v_{i}$ $=\frac{1}{2}(f_{i+1}+f_{i}),$
$u_{i}^{2}$ $\leq 2v_{i}w_{i}$
$f_{i}$ $=(1-\epsilon)\varphi(x_{i})+\epsilon\sum_{j=1}^{L}\varphi(x_{i}-\theta_{j})h_{j}$
$h$ $\in\mathcal{S}\equiv\{h\in\mathbb{R}^{L}:\sum_{j=1}^{L}h_{j}=1,\;h_{j}\geq 0,\;j=1,\dots,L\}$
Provided that the grids are sufficiently finely spaced interior point optimization in Mosek, ApS (2022), is capable of producing very accurate solutions very efficiently.
In Figure 7 we illustrate a Mallows marginal density, $f^{M}$, its corresponding mixing distribution, $P^{M}$, and plot the log mass of the discrete mass points of $P^{M}$ at their respective locations. On first glance, it seems that the mass points are approximately equally spaced and have mass that declines almost exponentially. However, on closer examination the spacing of the mass points in the right tail are estimated to be: $\{1.96,1.80,1.70,1.61,1.52,1.39,1.29,1.37,1.55,1.70,1.91\}$, which seems sufficiently non-uniform to call the uniform spacing conjecture into question. Donoho et al. (2013) suggest an alternative computational strategy for the Mallows problem using a parametric model that assumes equal spacing of the mass points. The grid for evaluation of $f$ is equally spaced from -30 to 30 with 500 points of evaluation. The grid for evaluation of $P^{M}$ is also equally spaced from -20 to 20 with 4003 points of evaluation. For purposes of illustration the mass at $\theta=0$ is taken to be 0.2.
Figure 7. Mallows least favorable marginal density, probability mass function of the Mallows mixing distribution and log mass of the mixing distribution as a function of location indicating the approximate exponentiality of the mixing distribution as conjectured by Mallows.
Figure 8. Mass points of the Mallows contamination distribution $H^{*}$ with $G_{0}=0.5\delta_{-2}+0.5\delta_{2}$. The solid black curve depicts the functional derivative of the objective function of the constrained Mallows’s problem. The vertical green lines depict the location of the mass points of the solution, while their length represents the amount of mass assigned to each. In accordance with the Tukey “hanging rootogram” principle these lengths are rescaled as the square root of the respective masses. The dashed horizontal line represents the bound on the pointwise risk imposed by the Hodges-Lehmann constraint in this case approximately 1.67 induced by the choice of $\epsilon=0.2$.
In the previous example we have taken the initial prior, $G_{0}$ to be point mass one at zero, but there is no obstacle to starting from any other initial prior. To provide some additional intuition about the nature of the Mallows solution for general $G_{0}$, we illustrate in Figure 8 a plot of the functional derivative of the objective evaluated at the solution,
$G^{*}(\theta)=(1-\epsilon)G_{0}(\theta)+\epsilon H^{*}(\theta)$
where $G_{0}(\theta)$ is taken to be an equally weighted mixture of two point masses at -2 and 2. The tangencies in this plot with the horizontal line $R(\delta^{*},\theta)=1+t$ coincide with the location of the mass points of the solution of $H^{*}$ indicated in the plot by the vertical green lines. The bound, $1+t$, is the pointwise risk bound chosen to constrain the modified prior, $G^{*}$ as determined by the choice of $\epsilon$. Since the initial prior places all its mass on the two points $\{-2,2\}$ the Hodges-Lehmann modification hedges this bet by placing a considerable mass at zero and exponentially declining mass at a few points below -2 and above +2. This figure is strongly reminiscent of a similar figure in Lindsay (1995) illustrating the location of mass points of the NPMLE.
Appendix B
Proof. [Proposition 3.1] For all $F\in\mathcal{F}_{\epsilon}$, we can represent its density as $f(x)=(1-\epsilon)f_{0}(x)+\epsilon h(x)$. When $f_{0}$ is symmetric around 0, which is our case here, then the optimal $f^{*}$ will necessarily have $h$ symmetric and hence $f^{*}$ itself symmetric around 0. Let $u(x)=\sqrt{f(x)}$, then $I(F)=4\int(u(x)^{\prime})^{2}dx$ with the constraint $u^{2}(x)\geq(1-\epsilon)f_{0}(x)$ and $\int u^{2}(x)dx=1$. The Lagrangean is then
$L(u)=4\int(u^{\prime}(x))^{2}dx+\lambda\Big(\int u(x)^{2}dx-1\Big)+\int\eta(x)\{(1-\epsilon)f_{0}(x)-u(x)^{2}\}dx.$
Denote $\psi(x)=\sqrt{(1-\epsilon)f_{0}(x)}$, then the KKT conditions for an optimal solution $u^{*}(x)=\sqrt{f^{*}(x)}$ are:
$(u^{*}(x))^{\prime\prime}$ $=\frac{\lambda-\eta(x)}{4}u^{*}(x)$
$\eta(x)(u^{*}(x)-\psi(x))$ $=0$
$u^{*}(x)$ $\geq\psi(x)$
$\eta(x)$ $\geq 0$
$\lambda$ $\geq 0.$
Define the two sets:
$\mathcal{C}$ $=\{x:u^{*}(x)=\psi(x)\}$
$\mathcal{F}$ $=\{x:u^{*}(x)>\psi(x)\}$
On set $\mathcal{F}$, $\eta(x)=0$, hence $(u^{*}(x))^{\prime\prime}=\frac{\lambda}{4}u^{*}(x)$. This differential equation has solution on each connected component of $\mathcal{F}$ of the form,
$u^{*}(x)=A\exp(kx/2)+B\exp(-kx/2),$
with $k=\sqrt{\lambda}$. For any closed subset of $\mathcal{F}$, say $|x|<c$ both exponential pieces are involved and $u^{*}(x)=A\cosh(kx/2)$. For any subset in the tail we need to kill one of the exponential terms to ensure integrability. Hence for the right tail component, with $x>b$, $u^{*}(x)=\sqrt{(1-\epsilon)f_{0}(b)}\exp(-k(x-b)/2)$ and for the left tail component with x<bx<-b, $u^{*}(x)=\sqrt{(1-\epsilon)f_{0}(-b)}\exp(k(x+b)/2)$. The remaining task is to find $(b,c,k,A)$ to guarantee that $f^{*}$ is in the feasible set $\mathcal{F}_{\epsilon}$. Continuity of $u^{*}(x)$ at the cut off points $\pm c$, implies $u^{*}(c)=\psi(c)$ which leads to,
(1) $A=\frac{\psi(c)}{\cosh(kc/2)}.$
We also require $(u^{*})^{\prime}(c)=\psi^{\prime}(c)$ and $(u^{*})^{\prime}(b)=\psi^{\prime}(b)$, These restrictions imply,
(2) $k=-(\log f_{0})^{\prime}(b)$
(3) $k\cdot\tanh(kc/2)=(\log f_{0})^{\prime}(c).$
Finally, $f^{*}$ has to integrate to 1, which implies,
(4) $\int_{-c}^{c}A^{2}\cosh^{2}(kx/2)+2\int_{c}^{b}(1-\epsilon)f_{0}(x)dx+\frac{2(1-\epsilon)f_{0}(b)}{k}=1.$
Solving (1) to (4) yields the constants, $(b,c,k,A)$.
Appendix C
In this appendix we demonstrate that the Bayes risk of the empirical Mallows decision rule, $\delta_{\hat{G}}^{M}$ converges that of the oracle Mallows rule, $\delta_{G}^{M}$, as $n\to\infty$ (Theorem 3.1). For notation convenience, we use $I(f)$ to refere to the Fisher information of the distribution $F$ whose density is $f$. That is, $I(f)=\int\frac{(f^{\prime})^{2}}{f}$. Let $f^{M}$ denote the minimizer of the Mallows problem:
$\min I(f)\quad\text{s.t.}f\in M,f\in\{\varphi*P,P\geq(1-\epsilon)G_{0}\}$
with $M:=\{\varphi*P:P\in\mathcal{P}(\mathbb{R})\}$. Let $\hat{f}^{M}_{n}$ be the minimizer of the empirical Bayes version of the Mallows problem
$\min I(f)\quad\text{s.t. }f\in M,f\in\{\varphi*P,P\geq(1-\epsilon)\hat{G}_{n}\}$
where $\hat{G}_{n}$ denotes the NPMLE of $G$. Denote $f_{0}=\varphi*G_{0}$ and $\hat{f}_{n}=\varphi*\hat{G}_{n}$. The Oracle robust decision rule is $\delta^{M}(x)=x+(\log f^{M}(x))^{\prime}$ and its empirical Bayes counter-part is $\hat{\delta}^{M}(x)=x+(\log\hat{f}_{n}^{M}(x))^{\prime}$. Denote the respective score functions by $g^{M}=(\log f^{M})^{\prime}$ and $\hat{g}^{M}_{n}=(\log\hat{f}_{n}^{M})^{\prime}$. We also know $f^{M}=\varphi*P^{M}$ and $\hat{f}^{M}_{n}=\varphi*P_{n}^{M}$ such that $P_{n}^{M}\in\{G:G=(1-\epsilon)\hat{G}_{n}+\epsilon H\}$ and $P^{M}\in\{G:G=(1-\epsilon)G_{0}+\epsilon H\}$.
The result of Theorem 3.1 requires the following assumption.
Assumption C.1.
As $n\to\infty$, $\hat{f}_{n}\to f_{0}$ strongly in $L^{2}$ and $I(f_{0})<\infty$.
Hellinger convergence is established in Jiang and Zhang (2009) under the assumptions they make on $G_{0}$. The strong convergence in $L^{2}$ is a consequence of Lemma C.6. The assumption $I(f_{0})<\infty$ satisfies automatically since $f_{0}\in M$. It provides a bound for $I(f^{M})$ since $f_{0}$ is in the feasible set of Mallows problem, hence $I(f^{M})\leq I(f_{0})<\infty$.
Proof. [Theorem 3.1] Suppose $X$ has density $f_{0}$, let the optimal Bayes rule under prior $G_{0}$ be $\delta^{B}(x)$. Denote the loss in Bayes risk of any decision rule $\delta$ in comparison to the optimal Bayes rule be,
$s(\delta,G_{0})=r(G_{0},\delta)-r(G_{0},\delta^{B}).$
We will show under Assumption C.1, as $n\to\infty$,
$s(\hat{\delta}^{M}(X),G_{0})-s(\delta^{M}(X),G_{0})\to 0.$
Consequently,
$r(G_{0},\hat{\delta}^{M})-r(G_{0},\delta^{M})\to 0.$
Let $g^{B}(x)=\delta^{B}(x)-x=(\log f_{0})^{\prime}$. By Stein’s lemma, we have
$s(\hat{\delta}_{n}^{M}(X),G_{0})-s(\delta^{M}(X),G_{0})$ $=\mathbb{E}_{f_{0}}[(\hat{g}_{n}^{M}(X)-g^{B}(X))^{2}]-\mathbb{E}_{f_{0}}[(g^{M}(X)-g^{B}(X))^{2}]$
$=\mathbb{E}_{f_{0}}[(\hat{g}_{n}^{M}-g^{M})(\hat{g}_{n}^{M}+g^{M}-2g^{B})].$
By Cauchy-Schwartz,
$|s(\hat{\delta}^{M}(X),G_{0})-s(\delta^{M}(X),G_{0})|\leq\sqrt{\mathbb{E}_{f_{0}}[(\hat{g}_{n}^{M}-g^{M})^{2}]}\sqrt{\mathbb{E}_{f_{0}}[(\hat{g}_{n}^{M}+g^{M}-2g^{B})^{2}]}.$
Theorem C.1 shows that the first term converges to 0 as $n\to\infty$. The second term
$\sqrt{\mathbb{E}_{f_{0}}[(\hat{g}_{n}^{M}+g^{M}-2g_{0})^{2}]}\leq\|\hat{g}_{n}^{M}\|_{L^{2}(f_{0})}+\|g^{M}\|_{L^{2}(f_{0})}+2\|g^{B}\|_{L^{2}(f_{0})}<C_{1}$
with $C_{1}$ a bounded constant. The last displayed result holds due to Lemma C.9 and the fact that $\|g^{B}\|_{L^{2}(f_{0})}=I(f_{0})<\infty$.
Theorem C.1.
Under Assumption C.1, we have as $n\to\infty$,
$\mathbb{E}_{f_{0}}[(\hat{g}^{M}_{n}-g^{M})^{2}]\to 0$
Proof. Consider the space
$\mathcal{T}:=\Big\{f\in L^{2}(\mathbb{R}):f\geq 0\ \text{a.e. and }\int_{\mathbb{R}}f(x)\,dx=1\Big\}.$
and we equip $\mathcal{T}$ with the subspace topology induced by the weak convergence of $L^{2}(\mathbb{R})$: a sequence $(f_{n})\subset\mathcal{T}$ converges to $f\in\mathcal{T}$ if and only if $f_{n}\rightharpoonup f$ in $L^{2}(\mathbb{R})$.
Fix $\epsilon\in(0,1)$ and define
$K_{n}:=\{\,f\in\mathcal{T}:\ \exists\,P\in\mathcal{P}(\mathbb{R})\ \text{with }f=\varphi*P\ \text{and}\ P\geq(1-\epsilon)\hat{G}_{n}\,\},$
$K:=\{\,f\in\mathcal{T}:\ \exists\,P\in\mathcal{P}(\mathbb{R})\ \text{with }f=\varphi*P\ \text{and}\ P\geq(1-\epsilon)G_{0}\,\}.$
Further define the functionals:
$F_{n}(f):=I(f)+\iota_{M}(f)+\iota_{K_{n}}(f),\qquad F(f):=I(f)+\iota_{M}(f)+\iota_{K}(f),$
where $\iota_{A}$ is the indicator function of the set $A$, taking value $\infty$ when violated.
The remainder of the proof is structured as follows:
  1. (1)
    We first establish that $F_{n}$ $\Gamma$-converge to $F$ in $\mathcal{T}$. (Lemma C.3) and $F_{n}$ is equi-coercive (Lemma C.2)
  2. (2)
    Then by Lemma C.1, the minimizer of $F$ is unique, we can invoke Corollary 7.24 in Dal Maso (1993) and obtain $\hat{f}^{M}_{n}\rightharpoonup f^{M}$ weakly in $L^{2}$ and $I(\hat{f}^{M}_{n})\to I(f^{M})$.
  3. (3)
    Using the fact that $\hat{f}^{M}\in M$ and $f^{M}\in M$, we further show $\hat{P}_{n}^{M}\Rightarrow P^{M}$ and $H(\hat{f}^{M}_{n},f^{M})\to 0$ as $n\to\infty$. (Theorem C.2)
Finally, Weak convergence $\hat{P}_{n}^{M}\Rightarrow P^{M}$ implies pointwise convergence for the density $\hat{f}^{M}_{n}$ and its derivative, hence
$\hat{g}_{n}^{M}(x)\to g^{M}(x),f_{0}-a.e.x$
Using the uniform bound (proven in Lemma C.9), we get
$|\hat{g}_{n}^{M}(x)-g^{M}(x)|^{2}\leq 4C^{2}<\infty$
so by dominated convergence,
$\int(\hat{g}_{n}^{M}-g^{M})^{2}f_{0}\to 0$
Lemma C.1.
Let $K:=\{\,f\in\mathcal{T}:\ \exists\,P\in\mathcal{P}(\mathbb{R})\ \text{with }f=\varphi*P\ \text{and}\ P\geq(1-\epsilon)G_{0}\,\}$, then $\min_{f\in M\cap K}I(f)$ has a unique minimizer.
Proof. $M$ is convex and $K$ is convex, hence $C=M\cap K$ is convex. Since $I(f)$ is strictly convex, the minimizer is unique.
Lemma C.2.
The sequence $(F_{n})$ is equi-coercive.
Proof. Fix $a\in\mathbb{R}$ and define the sublevel sets on the space $\mathcal{T}$
$E_{n,a}:=\{f\in\mathcal{T}:\ F_{n}(f)\leq a\},$
where $\mathcal{T}$ is equipped with the topology induced by weak convergence in $L^{2}$. By Definition 7.6 in Dal Maso (1993), it suffices to find, for each $a$, a set $K_{a}\subset\mathcal{T}$ that is compact in this topology and such that $E_{n,a}\subset K_{a}$ for all $n$.
Let $f\in E_{n,a}$. Then $F_{n}(f)<\infty$, hence $f\in M$. Thus $f=\varphi*P$ for some $P\in\mathcal{P}(\mathbb{R})$, and therefore
$0\leq f(x)=\int_{\mathbb{R}}\varphi(x-\theta)\,dP(\theta)\leq\|\varphi\|_{\infty}=\frac{1}{\sqrt{2\pi}}\qquad\forall x\in\mathbb{R}.$
Since also $\|f\|_{1}=\int f=1$, we obtain
$\|f\|_{2}^{2}\leq\|f\|_{\infty}\|f\|_{1}\leq\frac{1}{\sqrt{2\pi}},\qquad\text{i.e.}\qquad\|f\|_{2}\leq(2\pi)^{-1/4}.$
Hence, for every $n$ and every $a$,
$E_{n,a}\subseteq B:=\{f\in L^{2}(\mathbb{R}):\ \|f\|_{2}\leq(2\pi)^{-1/4}\}.$
Since $L^{2}(\mathbb{R})$ is reflexive, the closed ball $B$ is weakly compact. Therefore the set
$K_{a}:=\mathcal{T}\cap B$
is compact in $\mathcal{T}$ equipped with the subspace topology induced by weak convergence in $L^{2}$, and it satisfies $E_{n,a}\subseteq K_{a}$ for all $n$. This proves equi-coercivity.
Lemma C.3 ($\Gamma$-convergence in weak $L^{2}$).
Assume $\hat{f}_{n}\to f_{0}$ strongly in $L^{2}$. Then $F_{n}$ $\Gamma$-converges to $F$ in weak topology $L^{2}$.
Proof. We verify the $\Gamma$-liminf and $\Gamma$-limsup conditions with respect to weak $L^{2}$.
(i) $\Gamma$-liminf. Let $u_{n}\in\mathcal{T}$ and assume $u_{n}\rightharpoonup u$ in $L^{2}(\mathbb{R})$ for some $u\in\mathcal{T}$. We show $F(u)\leq\liminf_{n\to\infty}F_{n}(u_{n})$.
If $\liminf_{n\to\infty}F_{n}(u_{n})=+\infty$ there is nothing to prove. So we assume $\liminf_{n\to\infty}F_{n}(u_{n})<+\infty$. Set $\ell:=\liminf_{n\to\infty}F_{n}(u_{n})$. Choose indices $n_{k}\uparrow\infty$ such that $F_{n_{k}}(u_{n_{k}})\to\ell$ and $\sup_{k}F_{n_{k}}(u_{n_{k}})<\infty$. Since $u_{n_{k}}\rightharpoonup u$ and $u\in\mathcal{T}$, Lemma C.4 implies the measures $\mu_{k}(dx):=u_{n_{k}}(x)\,dx$ are tight.
For all large $k$, $F_{n_{k}}(u_{n_{k}})<\infty$ implies $u_{n_{k}}\in\mathcal{M}\cap K_{n_{k}}$. Hence there exist $P_{k}\in\mathcal{P}(\mathbb{R})$ such that $u_{n_{k}}=\varphi*P_{k}$ and $P_{k}\geq(1-\epsilon)\hat{G}_{n_{k}}$. By Lemma C.5 and tightness of $\{\mu_{k}\}$, the family $\{P_{k}\}$ is tight. By Prokhorov’s theorem, there exist $k_{j}\uparrow\infty$ and $P\in\mathcal{P}(\mathbb{R})$ such that $P_{k_{j}}\Rightarrow P$. Along the same sub-subsequence, since $\hat{G}_{n}\Rightarrow G_{0}$, $\hat{G}_{n_{k_{j}}}\Rightarrow G_{0}$.
Let $Q_{k_{j}}:=P_{k_{j}}-(1-\epsilon)\hat{G}_{n_{k_{j}}}\geq 0$. For any $\phi\in C_{b}(\mathbb{R})$ with $\phi\geq 0$,
$0\leq\int\phi\,dQ_{k_{j}}=\int\phi\,dP_{k_{j}}-(1-\epsilon)\int\phi\,d\hat{G}_{n_{k_{j}}}\longrightarrow\int\phi\,dP-(1-\epsilon)\int\phi\,dG_{0},$
so $P\geq(1-\epsilon)G_{0}$.
Next, for any $\psi\in C_{c}^{\infty}(\mathbb{R})\subset L^{2}(\mathbb{R})$,
$\int\psi\,u_{n_{k_{j}}}\to\int\psi\,u\qquad\text{and}\qquad\int\psi\,u_{n_{k_{j}}}=\int(\psi*\varphi)(\theta)\,dP_{k_{j}}(\theta)\to\int(\psi*\varphi)(\theta)\,dP(\theta)=\int\psi\,(\varphi*P).$
Hence $\int\psi\,u=\int\psi\,(\varphi*P)$ for all $\psi\in C_{c}^{\infty}(\mathbb{R})$, so $u=\varphi*P$ a.e. Therefore $u\in M$ and $u\in K$, hence $F(u)=I(u)$.
Finally, boundedness of $F_{n_{k_{j}}}(u_{n_{k_{j}}})$ implies $\sup_{j}I(u_{n_{k_{j}}})<\infty$. By weakly lower semi-continuity of $I(\cdot)$ (Lemma C.7)
$F(u)=I(u)\leq\liminf_{j\to\infty}I(u_{n_{k_{j}}})\leq\liminf_{j\to\infty}F_{n_{k_{j}}}(u_{n_{k_{j}}})=\ell,$
Hence concluding $F(u)\leq\ell=\liminf_{n\to\infty}F_{n}(u_{n})$.
(ii) $\Gamma$-limsup (recovery). Fix a $u\in\mathcal{T}$ such that $u=\varphi*G$ and $G\geq(1-\epsilon)G_{0}$. If $F(u)=+\infty$, take $u_{n}\equiv u$ and we are done. So we assume $F(u)<\infty$, hence $u\in M\cap K$. Define the non-negative finite measure $G_{v}:=G-(1-\epsilon)G_{0}\geq 0$, which has total mass $\epsilon$ and let $v:=\varphi*G_{v}$. Then we can rewrite $u=v+(1-\epsilon)f_{0}$ and define the recovery sequence $u_{n}=v+(1-\epsilon)\hat{f}_{n}$. Since $\hat{f}_{n}=\varphi*\hat{G}_{n}$, we can thus write
$u_{n}=\varphi*\Big(G_{v}+(1-\epsilon)\hat{G}_{n}\Big)$
which implies $u_{n}\in M\cap K_{n}\subset\mathcal{T}$. Since $\hat{f}_{n}\to f_{0}$ strongly in $L^{2}$ (Lemma C.6), hence $\hat{f}_{n}\rightharpoonup f_{0}$ in $L^{2}$, which implies $u_{n}\rightharpoonup u$ in $L^{2}$.
It remains to show $\limsup_{n\to\infty}F_{n}(u_{n})\leq F(u)$. Given our construction, $u\in M\cap K$ and $u_{n}\in M\cap K_{n}$ for every $n$, hence it suffices to show $\limsup_{n\to\infty}I(u_{n})\leq I(u)$. We will in fact show $I(u_{n})\to I(u)$, which is a stronger statement. To show that, since $\epsilon>0$, then $v(x)>0$ for all $x$. Defnote $\Psi(f):=(f^{\prime})^{2}/f$. Using inequality $\frac{(a+b)^{2}}{x+y}\leq\frac{a^{2}}{x}+\frac{b^{2}}{y}$ for $x,y>0$ with $a=v^{\prime}$, $b=(1-\epsilon)\hat{f}_{n}$, $x=v$, $y=(1-\epsilon)\hat{f}_{n}$, we obtain
$\Psi(u_{n})\leq\Psi(v)+(1-\epsilon)\Psi(\hat{f}_{n})$
We have $\hat{G}_{n}\Rightarrow G_{0}$, which implies $\hat{f}_{n}(x)\to f_{0}(x)$ pointwise and $\hat{f}_{n}^{\prime}(x)\to f_{0}^{\prime}(x)$ pointwise, hence $\Psi(u_{n}(x))\to\Psi(u(x))$ pointwise. Moreover apply Lemma C.8, we have
$\int\Big(\Psi(v)+(1-\epsilon)\Psi(\hat{f}_{n})\Big)=I(v)+(1-\epsilon)I(\hat{f}_{n})\to I(v)+(1-\epsilon)I(f_{0})$
We also have $I(v)<\infty$ since $v\in M$. Apply Pratt’s Lemma (C.10) we obtain $\int\Psi(u_{n})\to\int\Psi(u)$, e.g. $I(u_{n})\to I(u)$.
Theorem C.2.
Since $\hat{f}^{M}_{n}\in M$ and $f^{M}\in M$, let $\hat{f}^{M}_{n}:=\varphi*P_{n}^{M}\in M$ and $f^{M}:=\varphi*P^{M}\in M$ be densities. Assume $\hat{f}^{M}_{n}\rightharpoonup f^{M}$ in $L^{2}$, then $P_{n}^{M}\Rightarrow P^{M}$, $\hat{f}^{M}_{n}(x)\to f^{M}(x)$ for every $x$, and
$\|\hat{f}^{M}_{n}-f^{M}\|_{1}\to 0\qquad\text{and hence}\qquad H(\hat{f}^{M}_{n},f^{M})\to 0.$
Proof. Let $\varphi$ be density of $N(0,1)$ and set $\kappa_{x}(y):=\varphi(x-y)$. Since $\kappa_{x}\in L^{2}(\mathbb{R})$ and $\hat{f}^{M}_{n}\rightharpoonup f^{M}$ in $L^{2}$,
$h_{n}(x):=(\hat{f}^{M}_{n}*\varphi)(x)=\int\hat{f}^{M}_{n}(x)\,\varphi(x-y)\,dy=\langle\hat{f}^{M}_{n},\kappa_{x}\rangle\longrightarrow\langle f^{M},\kappa_{x}\rangle=:(f^{M}*\varphi)(x)=:h(x).$
Thus $h_{n}(x)\to h(x)$ for every $x$. Moreover, since $\hat{f}^{M}_{n}=\varphi*P_{n}^{M}$,
$h_{n}=\hat{f}^{M}_{n}*\varphi=(\varphi*P_{n}^{M})*\varphi=(\varphi*\varphi)*P_{n}^{M}=\psi*P_{n}^{M}.$
with $\psi$ being the density of $N(0,2)$.
Apply Lemma C.4, we have $(\hat{f}^{M}_{n}\,dx)$ is tight. Since $h_{n}:=\hat{f}_{n}^{M}*\varphi$, then $(h_{n}(x)\,dx)$ is tight by Lemma C.5.
Each $h_{n}$ is a density, so $\int h_{n}=1$. Also $0\leq h_{n}(x)\leq\|\psi\|_{\infty}$ for all $n,x$. By dominated convergence on [R,R][-R,R],
$\int_{-R}^{R}h_{n}(x)\,dx\to\int_{-R}^{R}h(x)\,dx.$
By tightness, choose $R$ with $\sup_{n}\int_{|x|>R}h_{n}\leq\eta$. Then
$1=\int h_{n}\leq\int_{-R}^{R}h_{n}+\eta.$
Letting $n\to\infty$ yields $1\leq\int_{-R}^{R}h+\eta\leq\int h+\eta$. Since also $\int h\leq\liminf_{n}\int h_{n}=1$ by Fatou, we conclude $\int h=1$. Therefore $h_{n}\to h$ pointwise a.e. with $\int h_{n}=\int h=1$, and Scheffé’s lemma implies
$\|h_{n}-h\|_{1}\to 0.$
For each $t\in\mathbb{R}$, let $\phi$ be the characteristic function,
$|\phi_{h_{n}}(t)-\phi_{h}(t)|=\left|\int e^{ity}(h_{n}-h)(y)\,dy\right|\leq\|h_{n}-h\|_{1}\to 0,$
so $\phi_{h_{n}}(t)\to\phi_{h}(t)$ for every $t$. But $h_{n}=\psi*P_{n}^{M}$, hence $\phi_{h_{n}}(t)=\phi_{\psi}(t)\phi_{P_{n}^{M}}(t)$. Since $\psi=N(0,2)$, $\phi_{\psi}(t)=e^{-t^{2}}\neq 0$, so
$\phi_{P_{n}^{M}}(t)=e^{t^{2}}\phi_{h_{n}}(t)\to e^{t^{2}}\phi_{h}(t)=:\phi_{P^{M}}(t).$
By Lévy’s continuity theorem, $P_{n}^{M}\Rightarrow P^{M}$.
For each fixed $x$, the map $\theta\mapsto\varphi(x-\theta)$ is bounded and continuous, so $P_{n}^{M}\Rightarrow P^{M}$ implies pointwise convergence:
$\hat{f}^{M}_{n}(x)=\int\varphi(x-\theta)\,dP^{M}_{n}(\theta)\to\int\varphi(x-\theta)\,dP^{M}(\theta)=f^{M}(x).$
Since $\hat{f}^{M}_{n},f^{M}$ are densities, apply Scheffé Lemma again gives $\|\hat{f}^{M}_{n}-f^{M}\|_{1}\to 0$, and then
$H^{2}(\hat{f}^{M}_{n},f^{M})=\int(\sqrt{\hat{f}^{M}_{n}}-\sqrt{f^{M}})^{2}\leq\int|\hat{f}^{M}_{n}-f^{M}|=\|\hat{f}^{M}_{n}-f^{M}\|_{1}\to 0.$
Lemma C.4 (Tightness from weak $L^{2}$ convergence to a density).
Let $(f_{n})_{n\geq 1}$ be probability densities on $\mathbb{R}$ such that:
  1. (1)
    $f_{n}\rightharpoonup f$ in $L^{2}(\mathbb{R})$ for some $f\in L^{2}(\mathbb{R})$;
  2. (2)
    $f$ is itself a probability density (i.e. $f\geq 0$ and $\int_{\mathbb{R}}f=1$);
Then the family of probability measures $(f_{n}(x)\,dx)$ is tight. More precisely: for every $\eta>0$ there exists $R<\infty$ such that
$\sup_{n\geq 1}\int_{|x|>R}f_{n}(x)\,dx\leq\eta.$
Proof. Fix $\eta>0$. Since $f$ is a probability density, choose $R_{0}$ such that
$\int_{|x|>R_{0}}f(x)\,dx\leq\eta/4.$
Let $\chi_{R_{0}}$ be a smooth cutoff function with
$0\leq\chi_{R_{0}}\leq 1,\qquad\chi_{R_{0}}\equiv 1\text{ on }[-R_{0},R_{0}],\qquad\operatorname{supp}(\chi_{R_{0}})\subset[-R_{0}-1,R_{0}+1].$
Then $\chi_{R_{0}}\in L^{2}(\mathbb{R})$, so weak $L^{2}$ convergence implies,
$\int_{\mathbb{R}}\chi_{R_{0}}(x)f_{n}(x)\,dx\longrightarrow\int_{\mathbb{R}}\chi_{R_{0}}(x)f(x)\,dx.$
Moreover,
$\int_{\mathbb{R}}\chi_{R_{0}}f\geq\int_{|x|\leq R_{0}}f(x)\,dx=1-\int_{|x|>R_{0}}f(x)\,dx\geq 1-\eta/4.$
Hence there exists $N$ such that for all $n\geq N$,
$\int_{\mathbb{R}}\chi_{R_{0}}(x)f_{n}(x)\,dx\geq 1-\eta/2.$
Since $\chi_{R_{0}}\leq\mathbf{1}_{\{|x|\leq R_{0}+1\}}$, we have
$\int_{|x|\leq R_{0}+1}f_{n}(x)\,dx\geq\int\chi_{R_{0}}f_{n}\geq 1-\eta/2,\qquad\text{so}\qquad\int_{|x|>R_{0}+1}f_{n}(x)\,dx\leq\eta/2\quad(n\geq N).$
For the finitely many indices $n=1,\dots,N-1$, choose $R_{1}$ large enough that
$\max_{1\leq n\leq N-1}\int_{|x|>R_{1}}f_{n}(x)\,dx\leq\eta/2.$
(Each individual $f_{n}$ is a probability density, so its tails go to 0 as $R\to\infty$.) Let $R:=\max\{R_{0}+1,R_{1}\}$. Then for $n\geq N$,
$\int_{|x|>R}f_{n}\leq\int_{|x|>R_{0}+1}f_{n}\leq\eta/2,$
and for $n\leq N-1$,
$\int_{|x|>R}f_{n}\leq\int_{|x|>R_{1}}f_{n}\leq\eta/2.$
Thus $\sup_{n}\int_{|x|>R}f_{n}\leq\eta$, proving tightness.
Lemma C.5.
Let $f_{n}$ be probability densities and define $h_{n}:=f_{n}*\varphi$, where $\varphi$ is the $N(0,1)$ density. If $(f_{n}(x)\,dx)$ is tight, then $(h_{n}(x)\,dx)$ is tight.
Proof. Let $X_{n}\sim f_{n}$ and let $Z\sim N(0,1)$ be independent. Then $X_{n}+Z$ has density $h_{n}$. For any $R>0$,
$\mathbb{P}(|X_{n}+Z|>R)\leq\mathbb{P}(|X_{n}|>R/2)+\mathbb{P}(|Z|>R/2).$
Given $\eta>0$, choose $R$ so that $\sup_{n}\mathbb{P}(|X_{n}|>R/2)\leq\eta/2$ (tightness of $f_{n}$) and $\mathbb{P}(|Z|>R/2)\leq\eta/2$. Then $\sup_{n}\mathbb{P}(|X_{n}+Z|>R)\leq\eta$, proving tightness of $(h_{n})$.
Lemma C.6.
For $\hat{f}_{n}\in M$ and $f_{0}\in M$, assume $H(\hat{f}_{n},f_{0})\to 0$ as $n\to\infty$, then,
$\hat{f}_{n}\to f_{0}\ \text{strongly in }L^{1}\quad\text{ and }\quad\hat{f}_{n}\to f\ \text{strongly in }L^{2}$
Proof. By Cauchy-Schwartz,
$\int|\hat{f}_{n}-f_{0}|\leq\Big(\int(\sqrt{\hat{f}_{n}}-\sqrt{f_{0}})^{2}\Big)^{1/2}\Big(\int(\sqrt{\hat{f}_{n}}+\sqrt{f_{0}})^{2}\Big)^{1/2}\leq 2H(\hat{f}_{n},f_{0})\to 0$
So we get the first result. Then note that,
$\int(\hat{f}_{n}-f_{0})^{2}=\int|\hat{f}_{n}-f_{0}|\cdot|\hat{f}_{n}-f_{0}|\leq\|\hat{f}_{n}-f_{0}\|_{\infty}\int|\hat{f}_{n}-f_{0}|$
Since $\hat{f}_{n}\in M,\|\hat{f}_{n}\|_{\infty}\leq\frac{1}{\sqrt{2\pi}}$ and the same for $f_{0}$,
$\|\hat{f}_{n}-f_{0}\|_{\infty}\leq\frac{2}{\sqrt{2\pi}}$
and combined with the established strong convergence in $L^{1}$, we get the strong convergence in $L^{2}$.
Lemma C.7.
Let $u_{n}\rightharpoonup u$ in $L^{2}$, then
$I(u)\leq\underset{n\to\infty}{\liminf}\ I(u_{n}).$
Proof. We need to show that Fisher information $I(\cdot)$ is weakly lower semi-continuous. Let $\phi\in C_{c}^{1}(\mathbb{R})$ be any continuously differentiable test function with compact support. Since $\int f(\frac{f^{\prime}}{f}+\phi)^{2}dx\geq 0$, expanding the quadratic, we have,
$\int\frac{(f^{\prime})^{2}}{f}+2\int f^{\prime}\phi+\int f\phi^{2}\geq 0.$
Note that the first term is $I(f)$, hence $I(f)\geq-2\int f^{\prime}\phi-\int f\phi^{2}$. Integrating by parts ($\int f^{\prime}\phi=-\int f\phi^{\prime}$), we have
$I(f)\geq 2\int f\phi^{\prime}-\int f\phi^{2}=\int f(2\phi^{\prime}-\phi^{2}),$
hence
$I(f)\geq\underset{\phi\in C_{c}^{1}(\mathbb{R})}{\sup}\int f(2\phi^{\prime}-\phi^{2}).$
On the other hand, pick $\eta_{R}\in C_{c}^{\infty}(\mathbb{R})$ with $0\leq\eta_{R}\leq 1$ and $\eta_{R}$ takes value 1 on [R,R][-R,R], and $\eta_{R}\uparrow 1$ pointwise as $R\to\infty$. Define
$\phi_{R}:=-\frac{f^{\prime}}{f}\eta_{R}\in C_{c}^{1}(\mathbb{R})$
Then using integration by parts and because $\phi_{R}$ is compactly supported,
$\int f(2\phi_{R}^{\prime}-\phi_{R}^{2})=-2\int f^{\prime}\phi_{R}-\int f\phi_{R}^{2}=2\int\frac{(f^{\prime})^{2}}{f}\eta_{R}-\int\frac{(f^{\prime})^{2}}{f}\eta_{R}^{2}=\int\frac{(f^{\prime})^{2}}{f}(2\eta_{R}-\eta_{R}^{2}).$
Since $2\eta_{R}-\eta_{R}^{2}\uparrow 1$ pointwise as $R\to\infty$, by monotone convergence theorem,
$\lim_{R\to\infty}\int f(2\phi_{R}^{\prime}-\phi_{R}^{2})=\int\frac{(f^{\prime})^{2}}{f}=I(f)$
hence $\sup_{\phi\in C_{c}^{1}(\mathbb{R})}\int f(2\phi^{\prime}-\phi^{2})\geq\sup_{R}\int f(2\phi_{R}^{\prime}-\phi_{R}^{2})\geq\lim_{R\to\infty}\int f(2\phi_{R}^{\prime}-\phi_{R}^{2})=I(f)$. Put together, we have
$I(f)=\sup_{\phi\in C_{c}^{1}(\mathbb{R})}\int f(2\phi^{\prime}-\phi^{2})$
For any fixed $\phi$, the map $f\to\int fg$ is weakly continuous in $L^{2}$ if $g\in L^{2}$. We have $g=2\phi^{\prime}-\phi$ is bounded and have compact support, hence
$f\to\int f(2\phi^{\prime}-\phi^{2})$
is weakly continuous in $L^{2}$. Then supremum of weakly continuous functionals is weakly lower semicontinuous, therefore if $u_{n}\rightharpoonup u$ in $L^{2}$,
$I(u)\leq\underset{n\to\infty}{\liminf}\ I(u_{n})$
Lemma C.8.
Suppose $\hat{f}_{n}\to f_{0}$ strongly in $L^{1}$, then
$I(\hat{f}_{n})\to I(f_{0}).$
Proof. Let $\psi$ be the characteristic function. We have the bound
$|\psi_{\hat{f}_{n}}(t)-\psi_{f_{0}}(t)|\leq\int_{\mathbb{R}}|\hat{f}_{n}(x)-f_{0}(x)|\,dx\to 0,\qquad\forall t\in\mathbb{R},$
hence $\psi_{\hat{f}_{n}}(t)\to\psi_{f_{0}}(t)$ pointwise on $\mathbb{R}$.
Since $\hat{f}_{n}=\varphi*\hat{G}_{n}$ and $f_{0}=\varphi*G_{0}$, their characteristic functions factorize as
$\psi_{\hat{f}_{n}}(t)=\psi_{\hat{G}_{n}}(t)\,e^{-t^{2}/2},\qquad\psi_{f_{0}}(t)=\psi_{G_{0}}(t)\,e^{-t^{2}/2}.$
Because $e^{-t^{2}/2}>0$ for all $t$, we obtain
$\psi_{\hat{G}_{n}}(t)=e^{t^{2}/2}\psi_{\hat{f}_{n}}(t)\to e^{t^{2}/2}\psi_{f_{0}}(t)=\psi_{G_{0}}(t)\qquad\forall t\in\mathbb{R}.$
By Lévy’s continuity theorem, $\hat{G}_{n}\Rightarrow G_{0}$.
For each fixed $x\in\mathbb{R}$, the maps $\theta\mapsto\varphi(x-\theta)$ and $\theta\mapsto\varphi^{\prime}(x-\theta)$ are bounded and continuous. Thus $\hat{G}_{n}\Rightarrow G_{0}$ (guaranteed since $\hat{G}_{n}$ is the NPMLE of $G_{0}$) implies
$\hat{f}_{n}(x)=\int\varphi(x-\theta)\,d\hat{G}_{n}(\theta)\to\int\varphi(x-\theta)\,dG_{0}(\theta)=f_{0}(x),$
$\hat{f}_{n}^{\prime}(x)=\int\varphi^{\prime}(x-\theta)\,d\hat{G}_{n}(\theta)\to\int\varphi^{\prime}(x-\theta)\,dG_{0}(\theta)=f_{0}^{\prime}(x).$
Since Gaussian mixtures are strictly positive everywhere, $f_{0}(x)>0$ and $\hat{f}_{n}(x)>0$ for all $x$, so the integrand
$\Psi(f)(x):=\frac{(f^{\prime}(x))^{2}}{f(x)}$
satisfies $\Psi(\hat{f}_{n})(x)\to\Psi(f_{0})(x)$ pointwise on $\mathbb{R}$.
Using $\varphi^{\prime}(t)=-(t)\varphi(t)$, we can write
$\hat{f}_{n}^{\prime}(x)=\int\varphi^{\prime}(x-\theta)\,d\hat{G}_{n}(\theta)=\int(\theta-x)\varphi(x-\theta)\,d\hat{G}_{n}(\theta).$
Applying Cauchy–Schwarz with respect to the probability measure $\hat{G}_{n}$ gives, for each $x$,
$\Psi(\hat{f}_{n})(x)$ $=\frac{\left(\int(\theta-x)\varphi(x-\theta)\,d\hat{G}_{n}(\theta)\right)^{2}}{\int\varphi(x-\theta)\,d\hat{G}_{n}(\theta)}$
$=\frac{\left(\int(\theta-x)\sqrt{\varphi(x-\theta)}\;\sqrt{\varphi(x-\theta)}\,d\hat{G}_{n}(\theta)\right)^{2}}{\int\varphi(x-\theta)\,d\hat{G}_{n}(\theta)}$
$\leq\frac{\left(\int(\theta-x)^{2}\varphi(x-\theta)\,d\hat{G}_{n}(\theta)\right)\left(\int\varphi(x-\theta)\,d\hat{G}_{n}(\theta)\right)}{\int\varphi(x-\theta)\,d\hat{G}_{n}(\theta)}$
$=\int(\theta-x)^{2}\varphi(x-\theta)\,d\hat{G}_{n}(\theta)=:D_{n}(x).$
Thus $0\leq\Psi(\hat{f}_{n})\leq D_{n}$ pointwise.
For each fixed $x$, the function $\theta\mapsto(\theta-x)^{2}\varphi(x-\theta)$ is bounded and continuous (boundedness holds since $\sup_{y\in\mathbb{R}}y^{2}\varphi(y)<\infty$). Hence $\hat{G}_{n}\Rightarrow G_{0}$ implies $D_{n}(x)\to D_{0}(x)$ pointwise, where
$D_{0}(x):=\int(\theta-x)^{2}\varphi(x-\theta)\,dG_{0}(\theta).$
By Fubini’s theorem and the change of variables $y=x-\theta$,
$\int_{\mathbb{R}}D_{n}(x)\,dx$ $=\int_{\mathbb{R}}\int_{\mathbb{R}}(x-\theta)^{2}\varphi(x-\theta)\,dx\,d\hat{G}_{n}(\theta)$
$=\int_{\mathbb{R}}\left(\int_{\mathbb{R}}y^{2}\varphi(y)\,dy\right)\,d\hat{G}_{n}(\theta)=\int_{\mathbb{R}}1\,d\hat{G}_{n}(\theta)=1.$
Similarly, $\int_{\mathbb{R}}D_{0}(x)\,dx=1$.
Taking stock, we have shown:
$0\leq\Psi(\hat{f}_{n})\leq D_{n},\qquad\Psi(\hat{f}_{n})\to\Psi(f_{0})\ \text{pointwise},\qquad D_{n}\to D_{0}\ \text{pointwise},\qquad\int D_{n}\to\int D_{0}=1.$
By Pratt’s lemma (Lemma C.10),
$\int_{\mathbb{R}}\Psi(\hat{f}_{n})(x)\,dx\to\int_{\mathbb{R}}\Psi(f_{0})(x)\,dx,$
i.e. $I(\hat{f}_{n})\to I(f_{0})$.
Lemma C.9.
Suppose $\hat{f}_{n}^{M}=\varphi*P^{M}_{n}\in M$ and $f^{M}=\varphi*P^{M}\in M$. Let $\hat{g}^{M}_{n}=(\log\hat{f}_{n}^{M})^{\prime}$ and $g^{M}=(\log f^{M})^{\prime}$ and define $\delta^{M}(x)=x+g^{M}(x)$ and $\hat{\delta}^{M}(x)=x+\hat{g}^{M}_{n}(x)$. Let $R(\theta,\delta)=\mathbb{E}_{Y\sim N(\theta,1)}[(\delta(Y)-\theta)^{2}]$. Assume $\sup_{n}\sup_{\theta\in\mathbb{R}}R(\theta,\hat{\delta}^{M})<\infty$ and $\sup_{\theta\in\mathbb{R}}R(\theta,\delta^{M})<\infty$, then there exists a universal constant $C$ such that
$\sup_{n}\|\hat{g}_{n}^{M}\|_{\infty}\leq C,\quad\|g^{M}\|_{\infty}\leq C$
Proof. Note that,
$R(\theta,\delta)=\mathbb{E}_{X\sim N(\theta,1)}[(\delta(X)-\theta)^{2}].$
If $\delta$ is nondecreasing, then $\delta(\theta)$ is a median of $\delta(X)$ when $X\sim N(\theta,1)$. This implies that,
$R(\theta,\delta)\geq\frac{(\delta(\theta)-\theta)^{2}}{2}.$
And,
$\mathbb{E}[(\delta(X)-\theta)^{2}]\geq\mathbb{E}[(\delta(X)-\theta)^{2}1\{\delta(X)\geq\delta(\theta)\}]\geq(\delta(\theta)-\theta)^{2}\mathbb{P}(\delta(X)\geq\delta(\theta))=\frac{1}{2}(\delta(\theta)-\theta)^{2}.$
Since $g(x)=\delta(x)-x$ for all $x$, let $x=\theta$, we have $g(\theta)=\delta(\theta)-\theta$, hence
$R(\theta,\delta)\geq(g(\theta))^{2}/2$
Note $\sup_{\theta}R(\theta,\delta)\leq 1+\bar{t}$, then
$\|g\|_{\infty}\leq\sqrt{2(1+\bar{t})}$
Lastly, $\delta^{M}$ and $\hat{\delta}^{M}$ are monotone function since by Tweedie formula, they are the posterior means under the priors $P$ and $P_{n}$ respectively. Applying the inequality to both $\delta^{M}$ and $\hat{\delta}^{M}$, we obtain the conclusion.
Lemma C.10 (Pratt’s lemma, Theorem 5.5 in Gut (2013)).
Let $X_{n},\,Y_{n},\,X,\,Y$ be nonnegative measurable functions on a measure space $(\Omega,\mathcal{A},\mu)$ such that $0\leq X_{n}\leq Y_{n}$ for all $n$, $X_{n}\to X$ a.e., $Y_{n}\to Y$ a.e., and $\int Y_{n}\,d\mu\to\int Y\,d\mu<\infty$. Then $\int X_{n}\,d\mu\to\int X\,d\mu$.